OpenStudy (amistre64):

If the equation \(A\vec x = \vec b\) has atleast one solution for each \(\vec b\) in \(R^n\), then the solution is un ique for each \(\vec b\). True or False?

OpenStudy (amistre64):

say \(A=\begin{pmatrix}1&0&3\\0&1&4\end{pmatrix} \) wouldnt this suggest that b has at least one solution, and that it is not unique?

OpenStudy (amistre64):

i forgot to read: "assume A is an nxn matrix ..." hmm but still it doesnt suggest that its linearly independant does it

OpenStudy (kinggeorge):

I think you also have to consider that it has at least one solution for every \(\vec b \in \mathbb{R}^n\). And that A is an \(n\times n\) matrix...

OpenStudy (amistre64):

2 solutions is "at least" one solution ...

OpenStudy (kinggeorge):

In that case, I'm pretty sure it's true. Especially since \(\mathbb{R}^n\) is a finite field extension over \(\mathbb{R}\).

OpenStudy (amistre64):

that sounds abstract albegra-y

OpenStudy (kinggeorge):

On the other hand, what if A doesn't have an inverse. Would that change things?

OpenStudy (amistre64):

So, if an Anxn hits every b in R^n; then it would span R^n .... right?

OpenStudy (amistre64):

if it spans then its linearly independant; has in inverse .. yada yada yada

OpenStudy (kinggeorge):

Right. So this statement would be true then.

OpenStudy (amistre64):

yeah, I just dint read all the pertinent information to begin with on the exam :) it makes more sense now